# Optimal Approximations for the Requirement Cut Problem on Sparse Graph Classes

### v2

@@ 2 Preliminaries
We begin by defining the key concepts used throughout the paper.
@@
For the two next definitions, let $G=(V,E)$ and $\mathcal{S}=\{S_{1},\ldots,S_{g}\}\subseteq V$ a set of terminal sets.
@@
A Steiner tree for a terminal set $S$ is a minimally connected subgraph $T\subseteq G$ that spans all vertices in $S$. $T$ may include additional non-terminal vertices, called Steiner vertices. An edge connecting two terminals (resp. Steiner) vertices may be called terminal edge (resp. Steiner edge).
@@
With this definition of Minimal Steiner tree, let us prove the following claim:
@@
In a graph $G=(V,E)$ with $\mathcal{S}=\{S_{1},\ldots,S_{g}\}$ terminal sets, we have : $\sigma(G,\mathcal{S})\leq g\cdot\tau(G)$.
@@
Let $T$ be a minimal Steiner tree of set $S_{k}$ for some $1\leq k\leq g$. By adding edges and vertices to $T$ until it covers $V(G)$ one can create at least one spanning tree per minimal Steiner tree. Also two distinct minimal Steiner trees (even of the same terminal set) can not extend to the same spanning tree as they differ in at least one edge.
Definition 2.6 (Series-Parallel Graph).
The class of series-parallel graphs is recursively defined as follows.
  • The graph consisting of a single edge $(x,y)$ is series-parallel with terminals $x$ and $y$.
  • Let $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ be two series-parallel graphs with terminals $x_{1},y_{1}\in V_{1}$ and $x_{2},y_{2}\in V_{2}$. Then:
    • The series composition of $G_{1}$ and $G_{2}$ identifies $y_{1}$ with $x_{2}$ (creating a new vertex $z$), and results in a new graph with terminals $x=x_{1}$ and $y=y_{2}$.
    • The parallel composition identifies $x_{1}$ with $x_{2}$ (creating $x$), and $y_{1}$ with $y_{2}$ (creating $y$), resulting in a new graph with terminals $x$ and $y$.
@@
Let $G$ be a series-parallel graph and let $T$ be its composition trace. The depth of $G$, is defined as the depth of the tree $T$, that is, if $l$ is the length of the longest path from a leaf (corresponding to a single edge) to the root (corresponding to $G$ itself): 0pt$=l-1$ Equivalently, it is the number of composition steps along the longest sequence of nested series and parallel compositions used to construct $G$.
For example, a single edge would be a series-parallel graph of depth 0, a path or a many parallel edges would have depth 1 and a composition of parallel paths would have depth 2.
@@ 3.1 Graphs with Polynomial number of Spanning Trees
A key parameter in our analysis is the number of spanning trees in a graph, which we denote as $\tau(G)$. By Kirchhoff’s Matrix Tree Theorem [30], $\tau(G)$ can be computed as the determinant of any cofactor of the graph Laplacian matrix of the graph. Equivalently, if $\lambda_{2},\lambda_{3},\dots,\lambda_{n}$ are the nonzero eigenvalues of $L$, then $\tau(G)=\frac{1}{n}\prod_{i=2}^{n}\lambda_{i}$. In our setting, we focus on graphs where $\tau(G)$ is polynomial in $n$, i.e., $\tau(G)=\mathcal{O}(n^{c})$ for some constant $c$.
The number of spanning trees multiplied by $g$ (the number of terminal sets of our Requirement Cut problem instance) is a natural upper bound to the number of distinct minimal Steiner trees in that same instance. So naturally, graphs with a polynomial number of spanning trees satisfy our condition. Let us now see three examples of graph classes that satisfy our spanning tree condition.
@@ 3.1.1 Graphs with Small Feedback Edge Sets
One prominent category of graphs with polynomially many spanning trees is the graphs that have a small, or bounded, feedback edge set. The feedback edge set is defined as the minimum number of edges whose removal results in an acyclic graph. Note that the feedback edge set number of a connected n-vertex and m-edge graph G is always $m-n+1$ and a minimum feedback edge set can be determined by the computation of a spanning tree in $\mathcal{O}(n+m)$ time via depth-first search. Also note that this class of graphs contains graphs with bounded number of cycles.
Back to our case, if the size of a graph’s feedback edge set is bounded by a constant $k$, then the number of spanning trees is $\binom{n-1+k}{k}=\mathcal{O}(n^{k})$, and hence polynomially bounded in the size of the graph.
3.1.2 Graphs with Short Cycles
@@ 3.2 Instances of Polynomial number of Steiner Trees
The Steiner trees in which we are interested for our algorithm are minimal Steiner trees, as given by Definition 2.4. We denote the number of such trees for a given instance by $\sigma(G,\mathcal{S})$. This value depends on the terminal sets of the instance as well as on the graph, and is hence more specific than its upper bound $\tau(G)\cdot g$. Unfortunately it is hard to compute. Still, it can provide better approximation for some type of instances where the $\tau(G)$ would yield a value that is super-polynomial in $n$.
Those instances are the instances where the whole graph may have a super-polynomial number of spanning trees but the number of Steiner trees is limited. This can occur for example if the terminal sets are each fully located in subgraphs that have a polynomial number of spanning trees when cut from the rest of the graph and that edge-cut to the rest of the graph is of size 1.
@@
Figure 4: Instance where terminal groups reside in minimally connected subgraphs. Here, it would be enough that $\tau(H_{i})=poly(|H_{i}|),i\in\{1,2,3\}$ to have a $\log(n)$ approximation for the instance.
@@ 4.1 Linear Program
We work with essentially the same $LPs$ as [32] and [25]. However, in Lemma 4.1, we will provide additional insights about these linear programs, compared to previous works. This is central to our analysis as it allows us to guarantee that our rounded solution is indeed a valid solution to the Requirement Cut problem. Both these $LPs$ work on slightly modified graphs. The first, IP-Steiner, calls for a complete graph, so we add 0-cost edges to the graph until it is a clique, we call this graph completed $G_{V}$. The second, Relaxed LP, calls for extra edges only between terminals of the same group. We shall call this modified graph, where $S_{i}$ is a clique with the addition of 0-cost edges, $G_{S_{i}}=(V,E_{S_{i}})$. This second modification is slightly different to the one used by [32] and [25] who used the induced graph $G[S_{i}]$ for the first and the complete graph on $S_{i}$ ($G=(S_{i},\binom{S_{i}}{2})$for the other. Our relaxation is equivalent when it comes to the purpose of this paper.
[…]
@@
(Relaxed $(\text{Relaxed LP})\quad\min$ $\sum_{e\in E}c_{e}d_{e}$
@@
IP-Steiner is an exact formulation of the Requirement Cut problem. However, we can notice that it is not clear why with only the constraint on the number of cut edges per Steiner tree, the solution would be a solution to the Requirement Cut problem. In Lemma 4.1 we will precisely show the role of the completeness of the graph in IP-Steiner. In fact, without the completeness of the graph in IP-Steiner (or of the terminal sets in Relaxed LP), Steiner trees could have a long $d$ length without satisfying the requirement. For example, take a path that starts and ends with terminal vertices and arbitrarily many Steiner vertices in the middle. This graph could be cut $r-1$ times satisfying the constraint on Steiner trees but having its terminal lying only in two components and not $r$.
The algorithm we present in this section is a randomized rounding of the Relaxed LP solution which can be solved by an ellipsoid method, using an MST algorithm as an oracle. What it ensures is that each and every Steiner tree is cut at least $r-1$ times. This alone does not suffice for proving that it returns a valid solution. Actually in [32] Nagarajan and Ravi use a preprocessing of their graph allowing them to have a relation between number of edges removed from the graph and number of terminal components created (Claim 2 in their paper). This same claim is used in Gupta, Nagarajan and Ravi [25] (Claim 3 in that paper) and it can be used simply because they are working with trees. In our case, we need something slightly stronger and more general since we are working with a more general class of graphs. Hence, before presenting our algorithm we will show an equivalence (under the condition of 0-cost edges added of our $LP$) between the two statements “All terminals of a set $S_{i}$ are in $r_{i}$ components” and “All spanning trees in $G_{S_{i}}$ are cut at least $r_{i}-1$ times.” with $G_{S_{i}}$ being the graph $G$ with extra 0-cost edges making the vertices of $S_{i}$ a clique. Also, just like the two papers mentioned, let $d^{*}$ be the solution to Relaxed LP, we work with $d_{\{u,v\}}=\min(2\cdot d^{*}_{\{u,v\}},1)$ which is also a metric since $d^{*}$ is one.
Claim 2 ([32]).
For any group $S_{i},i\in(1,\dots,g)$, the minimum Steiner tree on $S_{i}$ under metric $d$ has length at least $r_{i}-1$.
@@
Let $G=(V,E)$ a graph, $\mathcal{S}=\{S_{1},\ldots,S_{g}\}$ with $S_{i}\subseteq V$ a collection of terminal sets, and integer requirements $r_{1},\ldots,r_{g}\geq 2$. Let each $G_{S_{i}}$ be the graph where $S_{i}$ is connected by 0-cost edges making it a clique. Let $d:E\rightarrow\{0,1\}$ a metric such that for every $S_{i}$-Steiner tree $T$, $d(T)\geq r_{i}-1$. Then it is equivalent to say that all spanning trees over $G_{S_{i}}$ are cut at least $r_{i}-1$ times and to say that the vertices of $S_{i}$ lie in $r_{i}$ components.
@@
For the purpose of this proof we drop the indexes by considering the graph $G_{S}=(V,E_{S})$ with $S$ being a terminal set with requirement $r$. It is enough to show the property for one terminal set since more terminal sets only cut the graph in more components by cutting more edges. We call an edge $e$ with $d(e)=1$ a cut-edge. We call a component containing a terminal vertex, a terminal component. A non-terminal component is a component that does not contain a terminal vertex.
Let $T_{0}$ be an $S$-Steiner tree in $G$. $d(T_{0})\geq r-1$ and $S$ in less than $r$ components of $T_{0}=(k_{1},\dots,k_{\chi})$ with $\chi\geq r$. Let $k^{*}\in T_{0}$ be a non-terminal component. We know that $k^{*}$ is not a leaf component since $T_{0}$ is a minimal Steiner tree and hence it has at least 2 cut-edges incident to it, we note $c(k^{*})\geq 2$ this number. Also, let $t_{1},t_{2}\subseteq T_{0}\setminus k^{*}$ be two terminal subtrees adjacent to $k^{*}$. Let $x\in t_{1}$ and $y\in t_{2}$ two terminal vertices and $e^{*}$ the 0-cost edge in $G_{S}$ connecting them. Now we have $T_{1}$ another $S$-Steiner tree with $T_{1}=\{e^{*}\}\cup T_{0}\setminus k^{*}$. By our construction $d(T_{1})\geq d(T_{0})-c(k^{*})+1$ the $+1$ being $e^{*}$, by property of $d$, $d(T_{1})\geq r-1$. Now we have 2 case:
@@
    $d(T_{0})-c(k^{*})+1<r-1$: Then $T_{1}$ need to have a cut in one of the terminal subtrees effectively increasing its terminal components (along with $T_{0}$’s) by $1$ because it enters $t_{1}$ and $t_{2}$ at a terminal. We can then iterate over other non terminal components in $T_{0}\cup T_{1}$.
@@
    $d(T_{0})-c(k^{*})+1\geq r-1$: Then $T_{1}$ does not need anymore cut edges. In this case we have two subcases:
    1. (a)
      In $T_{1}$, $S$ lies in $r$ components: This contradicts the assumption since this would mean that $T_{0}$ lied in $r$ components.
    2. (b)
      In $T_{1}$, $S$ lies in less than $r$ components and so we have the same procedure creating $T_{2}$ that shortcuts one of the non-terminal components of $T_{1}$.
For case $(1)$ we can see that it increases the number of terminal components until we have created enough in some $T_{i}$ down the exploration.
For case $2b$ we have eliminated $k^{*}$ from $T_{1}$ and we can now attack and shortcut another non terminal component. This will necessarily converge since at each iteration we reduce by at least one the $d(T_{i})$.
This achieves the proof of the lemma.
Now that we proved Lemma 4.1, we can show our Algorithm that will ensure that each Steiner tree is cut with high probability enough times gives us a solution to the Requirement Cut problem with the same probability.

### v3

@@ 2 Preliminaries
Let $G=(V,E)$ and $\mathcal{S}=\{S_{1},\ldots,S_{g}\}\subseteq V$ a set of terminal sets. A Steiner tree for a terminal set $S$ is a minimally connected subgraph $T\subseteq G$ that spans all vertices in $S$ and is called an $S$-Steiner tree. $T$ may include additional non-terminal vertices, called Steiner vertices. An edge connecting two terminals (resp. Steiner) vertices may be called terminal edge (resp. Steiner edge).
Additionnaly we define the length of a Steiner tree $T$ as $\sum_{e\in T}d_{e}$ for a given length function $d$ and $d_{e}$ the length of edge $e$.
@@
With this definition of Minimal Steiner tree, let us prove the following claim: main-pratenddefaultcategory.tex
@@
In a graph $G=(V,E)$ with $\mathcal{S}=\{S_{1},\ldots,S_{g}\}$ terminal sets, we have $\sigma(G,\mathcal{S})\leq g\cdot\tau(G)$. With $\tau(G)$ the number of distinct spanning trees in $G$.
@@
Let $T$ be a minimal Steiner tree of set $S_{k}$ for some $1\leq k\leq g$. By adding edges and vertices to $T$ until it covers $V(G)$, one can create at least one spanning tree per minimal Steiner tree. Also, two distinct minimal Steiner trees for the same terminal set cannot extend to the same spanning tree in $G$, as they differ in at least one edge. Therefore, there exist at most $\tau(G)$ minimal Steiner trees for each terminal set $S_{i}\in\mathcal{S}$.
Definition 2.6 (Series-Parallel Graphs).
A series–parallel graph (SP graph) is any graph that can be obtained from a single edge $K_{2}$ by a finite sequence of the following operations:
  • Parallel composition: Replace an edge $(u,v)$ into a set of two parallel edges that connect $u$ and $v$.
  • Series composition: Replace an $(u,v)$ into a degree 2 vertex connected to $u$ and $v$.
@@
Let $G$ be a series-parallel graph and let $T$ be its composition trace. The depth of $G$, is defined as the depth of the tree $T$, that is, if $l$ is the length of the longest path from a leaf (corresponding to a single edge) to the root (corresponding to $G$ itself): 0pt=l1=l-1 .
For example, a single edge would be a series-parallel graph of depth 0, a path or a many parallel edges would have depth 1 and a composition of parallel paths would have depth 2. This means that this definition makes each class of bounded depth series-parallel graphs an infinite class of graphs.
@@ 3.1 Graphs with Polynomial number of Spanning Trees
A key parameter in our analysis is the number of spanning trees in a graph, which we denote as $\tau(G)$. By Kirchhoff’s Matrix Tree Theorem [kirchhoff_ueber_1847], $\tau(G)$ can be computed as the determinant of any cofactor of the graph Laplacian matrix of the graph. Equivalently, if $\lambda_{2},\lambda_{3},\dots,\lambda_{n}$ are the nonzero eigenvalues of $L$, then $\tau(G)=\frac{1}{n}\prod_{i=2}^{n}\lambda_{i}$. In our setting, we focus on graphs where $\tau(G)$ is polynomial in $n$, i.e., $\tau(G)=\mathcal{O}(n^{c})$ for some constant $c$.The number of spanning trees multiplied by $g$ (the number of terminal sets of our Requirement Cut problem instance) is a natural upper bound to the number of distinct minimal Steiner trees in that same instance. So naturally, graphs with a polynomial number of spanning trees satisfy our condition. First, let us now see two examples of graph classes that satisfy our spanning tree condition. Second we will also illustrate the kind of instances where the number of Steiner-tree is polynomial and for which any support graph, regardless of number of spanning trees, would admit a “single-log” approximation.
@@ 3.1.1 Graphs with Small Feedback Edge Sets
One prominent category of graphs with polynomially many spanning trees is the graphs that have a small, or bounded, feedback edge set. The feedback edge set is defined as the minimum number of edges whose removal results in an acyclic graph. Note that the feedback edge set number of a connected n-vertex and m-edge graph G is always mn+1m-n+1 and a minimum feedback edge set can be determined by the computation of a spanning tree in $\mathcal{O}(n+m)$ time via depth-first search. Also note that this class of graphs contains graphs with bounded number of cycles.Back to our case, if the size of a graph’s feedback edge set is bounded by a constant $k$, then the number of spanning trees is $\binom{n-1+k}{k}=\mathcal{O}(n^{k})$, and hence polynomially bounded in the size of the graph.
3.1.2 Graphs with few Short Cycles
@@ 3.2 Instances of Polynomial number of Steiner Trees
The Steiner trees we are interested in for our algorithm are minimal Steiner trees, as given by Definition 2.4. We denote the number of such trees for a given instance by $\sigma(G,\mathcal{S})$. This value depends on the terminal sets of the instance as well as on the graph, and is hence more specific than its upper bound $\tau(G)\cdot g$. Unfortunately, it is hard to compute. Still, it can provide better approximation for some type of instances where the $\tau(G)$ would yield a value that is super-polynomial in $n$.In those instances the whole graph may have a super-polynomial number of spanning trees but $\sigma(G,\mathcal{S})$ is in $poly(n)$. This can occur when the terminal sets are fully located in subgraphs that each have a polynomial number of spanning trees and are connected to the rest of the graph by exactly one edge.
@@
Figure 4: Instance where terminal groups reside in minimally connected subgraphs. Here, $\tau(H_{i})=poly(|H_{i}|),i\in\{1,2,3\}$ would suffice to have a $\log(n)$ approximation for the instance.
@@ 4.1 Linear Program
We work with essentially the same $LPs$ as [nagarajan_approximation_2010] and [gupta_improved_2010]. However, in Lemma 4.1, we will provide additional insights about these linear programs, compared to these previous works. This is central to our analysis as it allows us to guarantee that our rounded solution is indeed a valid solution to the Requirement Cut problem. Both these $LPs$ work on slightly modified graphs w.r.t $G=(V,E)$. The first $LP$, IP-Steiner, calls for a complete graph, so we add 0-cost edges to the graph until it is a clique, we call this graph $G_{V}$. The second, Relaxed LP, also calls for extra edges only between terminals of the same group. We shall call this modified graph, where $S_{i}$ is a clique with the addition of 0-cost edges, $G_{S_{i}}=(V,E_{S_{i}})$.
$(\text{IP-Steiner})\quad\min$ $\sum_{e\in E}c_{e}d_{e}$
@@
$(\text{Relaxed LP})\quad\min$ $\sum_{e\in E}c_{e}d_{e}$
@@
IP-Steiner is an exact formulation of the Requirement Cut problem. However, we can notice that it is not clear why with only the constraint on the number of cut edges per Steiner tree, the solution would be a solution to the Requirement Cut problem. In Lemma 4.1 we precisely show the role of the completeness of the graph in IP-Steiner. In fact, without the completeness of the graph in IP-Steiner (or of the terminal sets in Relaxed LP), Steiner trees could have a long $d$ length without satisfying the requirement. For example, take a path with any number of Steiner vertices in the middle and an arbitrary number of terminal vertices on each side. This graph could be cut r1r-1 times satisfying the constraint of the $LP$ while having its terminals lying only in 2 instead of $r$ components.The algorithm we present in this section is a randomized rounding of the Relaxed LP solution which can be solved by an ellipsoid method, using a Minimum Spanning Tree (MST) algorithm as an oracle. It ensures that each and every Steiner tree is cut at least r1r-1 times. This alone does not suffice for proving that it returns a valid solution. Actually in [nagarajan_approximation_2010, Claim 2] Nagarajan and Ravi use a preprocessing of their graph allowing them to have a relation between the number of edges removed from the graph and the number of terminal components created. This same claim is used in Gupta, Nagarajan and Ravi [gupta_improved_2010, Claim 3]. Both their approaches can use this claim because they work on trees. We need something stronger since we are working with a more general class of graphs. Hence, before presenting our algorithm we will show an equivalence (under the conditionof 0-cost edges added of our $LP$) between the two statements “All terminals of a set $S_{i}$ are in $r_{i}$ components” and “All spanning trees in $G_{S_{i}}$ are cut at least $r_{i}-1$ times.”. Also, just like the two papers mentioned, let $d^{*}$ be the solution to Relaxed LP, we work with $d_{\{u,v\}}=\min(2\cdot d^{*}_{\{u,v\}},1)$ which is also a metric since $d^{*}$ is one.
Claim 2 ([nagarajan_approximation_2010]).
For any group $S_{i},i\in(1,\dots,g)$, the minimum Steiner tree on $S_{i}$ under metric $d=\min(2\cdot d^{*},1)$ has length at least $r_{i}-1$.
@@
Let $G=(V,E)$ a graph, $\mathcal{S}=\{S_{1},\ldots,S_{g}\}$ with $S_{i}\subseteq V$ a collection of terminal sets, and integer requirements $r_{1},\ldots,r_{g}\geq 2$. Let each $G_{S_{i}}$ be the graph $G$ where 0-cost edges are added between nonadjacent pairs inside $S_{i}$ making $S_{i}$ a clique. Let $d:E\rightarrow\{0,1\}$ be a metric such that for every $S_{i}$-Steiner tree $T$, $d(T)\geq r_{i}-1$. Then it is equivalent to say that all spanning trees over $G_{S_{i}}$ are cut at least $r_{i}-1$ times and to say that the vertices of $S_{i}$ lie in $r_{i}$ components.
Summary of the proof: Starting from any $S$-Steiner tree whose cuts in $G$ do not split $S$ into $r$ terminal components, $T_{0}$. We successively find other trees that must have been considered by the LP and were also cut enough times. These other trees form a family of $S$-Steiner trees that use 0-cost edges and the union of their cut-edges in $G$ partitions $S$ into at least $r$ components.
@@
For the purpose of this proof we drop the indices by considering the graph $G_{S}=(V,E_{S})$ with $S$ being a terminal set with requirement $r$. It is enough to show the property for one terminal set since more terminal sets only cut the graph in more components by cutting more edges. We call an edge $e$ with $d(e)=1$ a cut-edge. We call a component containing a terminal vertex, a terminal component. A non-terminal component is a component that does not contain a terminal vertex.Let $T_{0}$ be an $S$-Steiner tree in $G$ such that $d(T_{0})=r-1$ (without loss of generality since more cut edges only work in our favor) and $S$ is in less than $r$ terminal components of $T_{0}=(K_{1},\dots,K_{r})$. Let $K^{*}\in T_{0}$ be a non-terminal component. Let $\delta(K^{*})$ be the number of cut-edges that have exactly one end-point in $^{*}. KK^{*}$ is not a leaf component since $T_{0}$ is a minimal Steiner tree and hence $K^{*}$ has at least 2 cut-edges incident to it, $\delta(K^{*})\geq 2$. Also, let $(t_{1},\dots,t_{\delta(K^{*})})\subseteq T_{0}\setminus K^{*}$ be the terminal subtrees (subtrees can contain several components) adjacent to $K^{*}$ through cut-edges $(e_{1},\dots,e_{\delta(K^{*})})$. Let $x\in t_{1}$ and $y\in t_{2}$ two terminal vertices and $e^{*}=(x,y)$ the 0-cost edge in $G_{S}$ connecting them. We call it the shortcut-edge.With all this, we can shortcut $K^{*}$ and reconstruct a tree $T_{1}$ that was considered by the LP and contains $e^{*}$ (the shortcut edge). We remove from the cut $e_{1}$ and $e_{2}$ the edges connecting $K^{*}$ to $t_{1}$ and $t_{2}$, and we add into the cut $e^{*}$ (necessary member of the cut since it costs 0 and achieves a desired separation) and an arbitrary edge of $T_{1}$, for example the closest to $e_{1}$ or $e_{2}$, say $e$. Now, we have $T_{1}$ another $S$-Steiner tree with $T_{1}=\{e^{*},e\}\cup T_{0}\setminus K^{*}$. By our construction so far $d(T_{1})=d(T_{0})=r-1$. More generally for any tree $T_{i}$ this procedure allows us to remove a non-terminal component and prove the existence of $T_{i+1}$ with a new set of edge-cuts.We effectively removed $K^{*}$ by short cutting it and we now have 2 cases:
@@
    The number of non-terminal components is reduced. If it is at 0, we are done since there are at least $r$ components.
@@
    The number of non-terminal components is the same.
Indeed, the number of non terminal components cannot increase and can at most create one. Case $(2)$ simply means that the new cut-edge $e$ cut a terminal component into a non-terminal part and a terminal part. However, in that case, we are sure to either get closer to decreasing the number of non-terminal components since we shrink the size of the terminal part of that terminal component or to actually decrease it that number.In some cases, the short cutting may shortcut more than one non-terminal components, in those cases, we reallocate all the cut edges of shortcut components (cut edges that are not in $T_{i+1}$) the same way we did for $e$ by shifting them to the nearest not cut edge of $T_{i+1}$.One last thing that remains to show is that we have enough shortcuts to shortcut all non-terminal components that may appear. Note that there at most r2r-2 non terminal components since at least $2$ components contain terminals. Also note that $r$ is bounded by the number of terminal vertices by definition of the Requirement Cut Problem. To shortcut those r2r-2 non terminal components, we have at least r1r-1 shortcuts since the worst case for number of shortcuts of available shortcut edges happens when many non-terminal components separate the same (up to a difference of other non-terminal components) terminal components. And the least number of shortcuts in that case happens when on one side of the non-terminal components there is exactly one terminal vertex and the rest on the other side. When this happens, we have r1r-1 shortcuts for r2r-2 non terminal components.
Figure 5: In this figure, the successive considered trees are in blue. The cut-edges indicated in green and the shortcut used to shortcut $K^{*}$ is in red like the terminal vertices. Notice how from the first to the second stage, we did not make progress in number of non-terminal components but still made progress towards cutting the terminals in the component.
Lemma 4.1 shows that it suffices for our algorithm to cut every Steiner tree enough times with high probability in order to yield a solution to the Requirement Cut problem.Note that even if the $LPs$ are solved on modified graphs with added 0-cost edges, the final cut is sampled on the original graph $G$ and all terminal sets are cut in $r$ components of $G$. Lemma 4.1 guarantees that cutting all Steiner trees in the modified graphs implies satisfying the requirements in $G$, therefore the validity and the cost of the solution remains intact and the number of Steiner trees remains the same since we can, by default, add all the 0 cost-edges to our cut if need be.